Calculate the heat stored Btu/ft2 in a concrete floor 6” thick if the initial temperature of the concrete is 60°F and the final temperature is 80°F
Answer #1

the density of concrete average = 140 lbs

mass of concrete = density * thickness *area

thickness= 6 inch = 0.5 ft

mass/ft2= 140 *0.5 =70 lbs /ft2

specific heat of concrete =0.2 Btu/lb- oF

heat stored = mass *specific heat * change in temperature

= 70 * 0.2*(80-60)

= 280 Btu/ft2

properties of concrete taken can be cleary seen in image:

PLEASE RATE THE SOLUTION BY LIKING IT...